jon has a piece of paper 22 inches long and 16 inches wide. He cuts a 2 inch by 2 inch square from each corner. He then folds up the sides of the paper and forms a box with an open top. What is the volume of his box?
Answer by mananth(16949)

jon has a piece of paper 22 inches long and 16 inches wide. He cuts a 2 inch by 2 inch square from each corner. He then folds up the sides of the paper and forms a box with an open top. What is the volume of his box?

jon has a piece of paper 22 inches long and 16 inches wide.
He cuts a 2 inch by 2 inch square from each corner.
length will be (22-4)=18
width will be (16-4)=12
The dimensions of the box will be 18 by 12 by 2
Volume = 18*12*2
=432 in^3